# Learning Constrained Markov Decision Processes With Non-stationary Rewards and Constraints

### v1

@@ F.3 Main results
We first decompose the regret as follows:
$R_{T}$ $=\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})$
$=\sum_{t\in[T]}(\mathbb{E}[r_{t}]-\overline{r})^{\top}(q_{t}-q^{*})+\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q^{*}-\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q_{t}$
$\leq\sum_{t\in[T]}\lVert\mathbb{E}[r_{t}]-\overline{r}\rVert_{1}+\sum_{t\in[T]}\left(\mathbb{E}[r_{t}]^{\top}q^{*}-\mathbb{E}[r_{t}]^{\top}q_{t}\right)$ (26a)
$\leq 2C+\sum_{t\in[T]}\left(\mathbb{E}[r_{t}]^{\top}q^{*}-\mathbb{E}[r_{t}]^{\top}q_{t}\right).$ (26b)
where Inequality (26a) holds since $|q_{t}(x,a)-q^{*}(x,a)|\leq 1,~{}\forall(x,a)\in X\times A$, and where Inequality (26b) holds by definition of $C$. Moreover, by Lemma 14, with probability at least $1-2\delta$, it holds:
$\sum_{t\in[T]}$ $\mathbb{E}[r_{t}]^{\top}q^{*}-\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q_{t}$
$=\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}\left(q^{*}-q_{t}^{j^{*}}\right)+\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}\left(q_{t}^{j^{*}}-q_{t}^{j_{t}}\right)$
$\leq R_{T}^{j^{*}}+\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}\mathbb{E}[r_{t}]^{\top}q_{t}^{j}-\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q_{t}^{j_{t}}-\frac{Lm+1}{\rho}\widehat{V}_{T}+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}}+$
@@
$\mkern 180.0mu+\mathcal{O}\left(\frac{M\ln T}{\eta}+\eta~{}{m^{4}ML^{4}T}+\gamma TLM+\frac{L}{\gamma}\ln\left(\frac{1}{\delta}\right)\right),$ (27)
@@
where Equation (27) holds by Lemma 14, with probability at least $1-2\delta$, and where $R_{T}^{j^{*}}$ is the regret $\texttt{Alg}^{j^{*}}$ would incur in. Now, employing the Azuma-Hoeffding inequality to $\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}\mathbb{E}[r_{t}]^{\top}q_{t}^{j}-\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q_{t}^{i_{t}}$ , Theorem 3 to bound $R_{T}^{j^{*}}$ with high probability considering that $2^{j^{*}}$ is at most $2C$ and Lemma 16 to bound $\widehat{V}_{T,j^{*}}$ with high probability, we observe that with probability at least $1-24\delta$, it holds:
$R_{T}^{j^{*}}$ $+\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}\mathbb{E}[r_{t}]^{\top}q_{t}^{j}-\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q_{t}^{j_{t}}+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}}$
@@
$\leq\mathcal{O}\left(m^{2}L^{2}|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m^{2}L\ln(T)|X||A|C+L^{2}m\frac{\ln(\frac{M}{\delta})}{\gamma}\right).$ (28)
@@
Finally, combining the previous results and by Union Bound, with probability at least $1-26\delta$, it holds:
@@
$R_{T}+\frac{Lm+1}{\rho}\widehat{V}_{T}\leq\mathcal{O}\Bigg{(}m^{2}L^{2}$ $|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m^{2}L\ln(T)|X||A|C+$
$+L^{2}m\frac{\ln(\frac{M}{\delta})}{\gamma}+\frac{M\ln T}{\eta}+\eta m^{4}ML^{4}T+\gamma TLM\Bigg{)},$ (29)
@@
which concludes the proof after observing that $\widehat{V}_{T}\geq 0$, by definition, and setting $\gamma=\sqrt{\frac{\ln(M/\delta)}{TM}}$, $\eta=\sqrt{\frac{\ln(T)}{T}}\frac{\rho}{2Lm(Lm+1)}$. ∎
@@
By Equation (29), which holds with probability at least $1-26\delta$, we obtain:
$R_{T}+\frac{Lm+1}{\rho}\widehat{V}_{T}\leq\mathcal{O}\Bigg{(}m^{2}L^{2}$ $|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m^{2}L\ln(T)|X||A|C+$
$+L^{2}m\frac{\ln(\frac{M}{\delta})}{\gamma}+\frac{M\ln T}{\eta}+\eta~{}m^{4}ML^{4}T+\gamma TLM\Bigg{)}.$
In order to obtain the final violations bound, it is necessary to find an upper bound for $-R_{T}$. We proceed as follows,
@@
$\leq\sum_{t\in[T]}\left(\sum_{i\in[m]}\left[\mathbb{E}[g_{t,i}]^{\top}q_{t}-\alpha_{i}\right]^{+}+\sum_{i\in[m]}\left(\left\lVert\overline{G}_{i}-g^{\circ}\right\rVert_{1}+\left\lVert g^{\circ}-\mathbb{E}[g_{t,i}]\right\rVert_{1}\right)\right)$
@@
$\leq\mathcal{O}\Bigg{(}m^{2}L^{2}|X||A|\sqrt{T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m^{2}L\ln(T)|X||A|C+$
$\mkern 180.0mu+L^{2}m\frac{\ln(\frac{M}{\delta})}{\gamma}+\frac{M\ln(T)}{\eta}+\eta m^{4}ML^{4}T+\gamma mTL^{2}M\Bigg{)},$
where the last inequality holds by Equation (29) and by Lemma 15, with probability at least $1-4\delta$ under $\mathcal{E}_{\widehat{q}}$. Employing a Union Bound, setting $\gamma=\sqrt{\frac{\ln(M/\delta)}{TM}}$ and $\eta=\sqrt{\frac{\ln(T)}{T}}\frac{\rho}{2Lijm(Lm+1)}$ concludes the proof. ∎

### v2

@@ F.3 Main results
Employing algorithm 3, with probability at least $1-14\delta$, it holds:
$R_{T}$ $=\sum_{t\in[T]}\overline{r}^{\top}q^{*}-\sum_{t\in[T]}\overline{r}^{\top}q_{t}$
$=\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t}^{j^{*}})+\sum_{t\in[T]}\overline{r}^{\top}(q_{t}^{j^{*}}-q_{t})$
$=\sqrt{\beta_{1}T}\nu_{T,j^{*}}+\beta_{2}\nu_{T,j^{*}}+2\beta_{3}C+\sum_{t\in[T]}\overline{r}^{\top}(q_{t}^{j^{*}}-q_{t})$ (28a)
$\leq\sqrt{\beta_{1}T}\nu_{T,j^{*}}+\beta_{2}\nu_{T,j^{*}}+2\beta_{3}C+2C-\frac{Lm+1}{\rho}\widehat{V}_{T}+\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}}$
$\mkern 50.0mu-(\sqrt{\beta_{1}}+\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}})\sqrt{T}\nu_{T,j*}-\left(\beta_{2}+\frac{m(mL+1)}{\rho}\beta_{5}\right)\nu_{T,j^{*}}$
$\mkern 50.0mu+\mathcal{O}\bigg{(}\frac{M\ln T}{\eta}+\eta~{}{m^{4}L^{4}TM}+\eta~{}M\ln(T)m^{4}L^{2}\left(\beta_{2}^{2}+\beta_{5}^{2}\right)$
@@
$\mkern 50.0mu+\eta T(\beta_{1}+L^{2}m^{4}\beta_{4})M\log(T)+\gamma TLM+L\sqrt{T\ln\left(\nicefrac{{1}}{{\delta}}\right)}+\frac{Lm}{\gamma}\ln\left(\nicefrac{{1}}{{\delta}}\right)\bigg{)}.$ (28b)
where Inequality (28a) hold with probability at least $1-11\delta$ by Corollary 7,Inequality (28b) holds with probability at least $1-3\delta$ thanks to Lemma 14 and to the following reasoning, which holds with probability at least $1-\delta$:
$\sum_{t\in[T]}\overline{r}^{\top}(q_{t}^{j*}-q_{t})$ $=\sum_{t\in[T]}(\overline{r}-\mathbb{E}[r_{t}])^{\top}(q_{t}^{j*}-q_{t})+\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}(q_{t}^{j*}-q_{t})$
$\leq\sum_{t\in[T]}\lVert\overline{r}-\mathbb{E}[r_{t}]\rVert_{1}+\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}\left(q_{t}^{j*}-q_{t}\right)$ (29a)
$\leq 2C+\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}\left(q_{t}^{j*}-q_{t}\right)$ (29b)
$\leq 2C+\sum_{t\in[T]}\mathbb{E}[r_{t}]^{\top}q_{t}^{j*}-\sum_{t\in[T]}\sum_{j\in[M]}w_{t,j}\mathbb{E}[r_{t}]^{\top}q_{t}^{j}+L\sqrt{2T\ln(\nicefrac{{1}}{{\delta}})}$ (29c)
where Inequality (29a) holds since $|q_{t}(x,a)-q_{t}^{j^{*}}(x,a)|\leq 1,~{}\forall(x,a)\in X\times A$, where Inequality (29b) holds by definition of $C$, and where Inequality (29c) use Azuma-Hoeffding inequality.
@@
We can apply Lemma 16 to bound $\widehat{V}_{T,j^{*}}$ with high probability. In fact we observe that with probability at least $1-16\delta$, it holds:
$\frac{Lm+1}{\rho}\widehat{V}_{T,j^{*}}$
@@
$\leq\mathcal{O}\left(m^{2}L^{2}|X|\sqrt{|A|T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m^{2}L\beta_{6}C+m^{2}L\ln(T)|X||A|C+L^{2}m^{2}\frac{\ln\left(\frac{M}{\delta}\right)}{2\gamma}\right)$
$\mkern 200.0mu+\frac{(Lm+1)m}{\rho}\beta_{5}\nu_{T,j^{*}}+\frac{m(Lm+1)}{\rho}\sqrt{\beta_{4}T}\nu_{T,j^{*}}.$
@@
Finally, combining the previous results and by Union Bound, with probability at least $1-30\delta$, it holds:
$R_{T}+\frac{Lm+1}{\rho}\widehat{V}_{T}$
$\leq\mathcal{O}\bigg{(}\frac{M\ln T}{\eta}+\eta~{}{m^{4}L^{4}TM}+\eta~{}M\ln(T)m^{4}L^{2}(\beta_{2}^{2}+\beta_{5}^{2})+\eta T(\beta_{1}+L^{2}m^{4}\beta_{4})M\log(T)$
@@
$\quad+\gamma TLM+L\sqrt{T\ln\left(\nicefrac{{1}}{{\delta}}\right)}+\frac{Lm}{\gamma}\ln\left(\nicefrac{{1}}{{\delta}}\right)$
$\quad+m^{2}L^{2}|X|\sqrt{|A|T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+mL\beta_{6}C+\beta_{3}C+m^{2}L|X||A|\ln(T)C\bigg{)}$ (30)
@@
which concludes the proof after observing that $\widehat{V}_{T}\geq 0$, by definition, and setting $\gamma=\sqrt{\frac{\ln(M/\delta)}{TM}}$, $\eta\leq\frac{1}{2\Lambda m\left(\sqrt{\beta_{1}T}+\beta_{2}+\beta_{5}+\sqrt{\beta_{4}T}\right)}$. ∎
@@
Starting from Inequality (30), in order to obtain the final violations bound, it is necessary to find an upper bound for $-R_{T}$. We proceed as follows,
@@
$\leq\sum_{t\in[T]}\left(\sum_{i\in[m]}\left[\mathbb{E}[g_{t,i}]^{\top}q_{t}-\alpha_{i}\right]^{+}+\sum_{i\in[m]}\left(\left\lVert\overline{G}_{i}-g^{\circ}_{i}\right\rVert_{1}+\left\lVert g^{\circ}_{i}-\mathbb{E}[g_{t,i}]\right\rVert_{1}\right)\right)$
@@
$\leq\mathcal{O}\Bigg{(}m^{2}L^{2}|X|\sqrt{|A|T\ln\left(\frac{mMT|X||A|}{\delta}\right)}+m^{2}L\ln(T)|X||A|C+\gamma mTL^{2}M\Bigg{)}$
$\mkern 400.0mu+\mathcal{O}\left(R_{T}+\frac{Lm+1}{\rho}\widehat{V}_{T}\right),$
where the last inequality holds by Equation (30) and by Lemma 15, with probability at least $1-4\delta$ under $\mathcal{E}_{\widehat{q}}$. Employing a Union Bound, setting $\gamma=\sqrt{\frac{\ln(M/\delta)}{TM}}$ and $\eta\leq\frac{1}{2\Lambda m\left(\sqrt{\beta_{1}T}+\beta_{2}+\beta_{5}+\sqrt{\beta_{4}T}\right)}$ concludes the proof. ∎
@@ G.3 Auxiliary lemmas for the transitions estimation
Appendix H Auxiliary lemmas for stability
In this section we state the results related to the stability of the arm-algorithms when $C$ is not known. The procedure is inspired by Jin et al. (2024) and Agarwal et al. (2017), but adapted to the case of Constrained MDP in high probability. We first give some important definitions. In these definitions we will use $C_{t}$ as the value of adversarial corruption at episode $t\in[T]$, where $C_{t}$ is defined as $C_{t}:=\max\{C_{t}^{G},C_{t}^{r}\}$, which meets the requirement of upper bounding the adversarial corruption at each considered episode. In addition it holds that $\sum_{t\in[T]}C_{t}\leq C_{r}+C_{G}$ or equivalently $C\leq\sum_{t\in[T]}C_{t}\leq 2C$, which does not influence the order of the analysis.
Definition 2.
A CMDP algorithm is corruption-robust if it takes $\theta$ (a guess on the corruption amount) as input, and achieves for any random stopping time $t^{\prime}\leq T$, whenever $\sum_{t\in[t^{\prime}]}C_{t}<\theta$:
$\sum_{t\in[t^{\prime}]}\overline{r}^{\top}(q^{*}-q_{t})\leq\sqrt{\beta_{1}t^{\prime}}+\left(\beta_{2}+\beta_{3}\theta\right)\mathbb{I}(t^{\prime}\geq 1),$
and
$\max_{i\in[m]}\sum_{t\in[t^{\prime}]}\left[g_{t,i}^{\top}q_{t}-\alpha_{i}\right]^{+}\leq\sqrt{\beta_{4}t^{\prime}}+\left(\beta_{5}+\beta_{6}\theta\right)\mathbb{I}(t^{\prime}\geq 1).$
Notice that Algorithm 2 is corruption-robust after applying a doubling trick to make it work for any stopping time, with probability at least $1-9\delta$ thank to Theorem 7 and Theorem 9 Furthermore, we introduce the notion of $\alpha$-stability. An algorithm is considered to be $\alpha$-stable, if its regret under condition imposed by Algorithm 3 is of order $\nu_{T}^{\alpha}\cdot\tilde{\mathcal{O}}\left(R_{T}\right)$, where $R_{T}$ is the upper bound on the regret attained by the algorithm if it receives feedback at each episode. In particular, we are interested in the $1$-stability.
Definition 3.
An algorithm is $1$-stable if, under the condition imposed by Algorithm 3, it holds:
$\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})\leq\sqrt{\beta_{1}T}\nu_{j,T}+\beta_{2}\nu_{j,T}+\beta_{3}C,$
and
$\max_{i\in[m]}\sum_{t\in[T]}\left[g_{t,i}^{\top}q_{t}-\alpha_{i}\right]^{+}\leq\sqrt{\beta_{4}T}\nu_{j,T}+\beta_{5}\nu_{j,T}+\beta_{6}C.$
We can use the procedure defined by Algorithm 4 and originally proposed by Jin et al. (2024) to transform a generic corruption robust algorithm to a $1$-stable algorithm. Differently from Jin et al. (2024), in our setting, we use the natural symmetry between regret and positive cumulative constraints violation to stabilize both the regret and the positive cumulative constraints violation. We have a different bound for $C_{t}$ (value of adversarial corruption at episode $t$): indeed, $C_{t}\leq\max\{\lVert\mathbb{E}[r_{t}]-r^{\circ}\rVert_{1},\max_{i\in[m]}\lVert\mathbb{E}[g_{t,i}]-g_{i}^{\circ}\rVert_{1}\}$ is bounded by $|X||A|$. Finally, we are interested in obtaining results that hold in high probability rather than in expectation. To do so, we focus on $1$-stability guarantee rather than $1/2$-stability as in Jin et al. (2024) since removing the expectation prevents us from achieving the result above with lower coefficients. We can state the following result.
Lemma 21 (Adapted from Jin et al. (2024)).
Given an algorithm which is corruption robust according to Definition 2 with parameters $(\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5},\beta_{6})$ and $\beta_{1}\geq\mathcal{O}(L^{2}\log(\nicefrac{{T}}{{\delta}}))$, $\beta_{4}\geq\mathcal{O}(L^{2}\log(\nicefrac{{T}}{{\delta}}))$, with probability at least $1-p$ with $p\in(0,1)$, then, it is possible convert it to an $1$-stable algorithm with probability at least $1-p-2\delta$ according to Definition 3 with parameters $(\beta_{1}^{\prime},\beta_{2}^{\prime},\beta_{3}^{\prime},\beta_{4}^{\prime},\beta_{5}^{\prime},\beta_{6}^{\prime})$ as $\beta_{1}^{\prime}=\mathcal{O}\left(\beta_{1}\right),\beta_{2}^{\prime}=\mathcal{O}\left(\beta_{2}+\beta_{3}|X||A|\log(\nicefrac{{\log(T)}}{{\delta}})\right),\beta_{3}^{\prime}=\mathcal{O}\left(\beta_{3}\log(T)\right),\beta_{4}^{\prime}=\mathcal{O}\left(\beta_{4}\right),\beta_{5}^{\prime}=\mathcal{O}\left(\beta_{5}+\beta_{6}|X||A|\log(\nicefrac{{\log(T)}}{{\delta}})\right),\beta_{6}^{\prime}=\mathcal{O}\left(\beta_{6}\log(T)\right)$, employing Algorithm 4.
Proof.
Suppose Algorithm 4 is initialized with the true value of adversarial corruption $C$. We will first prove the result for the regret. We will start by considering a generic instance algorithm $k\in[M]$. Define the quantity $d_{t,k}=\mathbb{I}(w_{t}\in(2^{-k-1},2^{-k}])$ and $h_{t,k}=\mathbb{I}(\text{Instance $k$ receives feedback at episode $t$})$. We observe that with probability at least $1-\left(p+\mathbb{P}\left(\bigcup_{k\in[\log(T)]}\{\sum_{t\in[T]}C_{t}d_{t,k}h_{t,k}>\theta_{k}\}\right)\right)$ it holds:
$\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})d_{t,k}h_{t,k}\leq\sqrt{\beta_{1}\sum_{t\in[T]}d_{t,k}h_{t,k}}+\left(\beta_{2}+\beta_{3}\theta\right)\max_{t\in[T]}d_{t,k},$
by the corruption-robust property of instance $k$. We study now the quantity $\mathbb{P}\left(\bigcup_{k\in[M]}\{\sum_{t\in[T]}C_{t}d_{t,k}h_{t,k}>\theta_{k}\}\right)$. Notice that $\mathbb{E}[h_{t,k}|d_{t,k}]=2^{-k-1}d_{t,k}$, and since $d_{t,k}$ is an indicator function then $\mathbb{E}[h_{t,k}|d_{t,k}]d_{t,k}=\mathbb{E}[h_{t,k}|d_{t,k}]$. In addition, since $\sum_{t\in[T]}C_{t}\leq 2C$, it holds:
$\sum_{t\in[T]}C_{t}\mathbb{E}[h_{t,k}|d_{t,k}]d_{t,k}=2^{-k-1}\sum_{t\in[T]}C_{t}d_{t,k}\leq 2^{-k}C,$
and with probability at least $1-\nicefrac{{\delta}}{{\log(T)}}$ noticing that $M=\log(T)$:
$\sum_{t\in[T]}C_{t}$ $d_{t,k}h_{t,k}-\sum_{t\in[T]}C_{t}\mathbb{E}[h_{t,k}|d_{t,k}]d_{t,k}$
$\leq 2\sqrt{\sum_{t\in[T]}C_{t}^{2}d_{t,k}\mathbb{E}[h_{t,k}|d_{t,k}]\log\left(\frac{\log(T)}{\delta}\right)}+|X||A|\log\left(\frac{\log(T)}{\delta}\right)$ (34a)
$\leq 2\sqrt{|X||A|\sum_{t\in[T]}C_{t}d_{t,k}\mathbb{E}[h_{t,k}|d_{t,k}]\log\left(\frac{\log(T)}{\delta}\right)}+|X||A|\log\left(\frac{\log(T)}{\delta}\right)$ (34b)
$\leq\sum_{t\in[T]}C_{t}\mathbb{E}[h_{t,k}|d_{t,k}]d_{t,k}+2|X||A|\log\left(\frac{\log(T)}{\delta}\right),$ (34c)
where Inequality (34a) holds with probability at least $1-\nicefrac{{\delta}}{{\log(T)}}$ by Freedman inequality, Inequality (34b) holds since $C_{t}\leq|X||A|$, and Inequality (34c) holds by AM-GM inequality. Therefore, it holds simultaneously for all $k\in[M]$:
$\sum_{t\in[T]}C_{t}d_{t,k}h_{t,k}$ $\leq 2\sum_{t\in[T]}C_{t}\mathbb{E}[h_{t,k}|d_{t,k}]d_{t,k}+2|X||A|\log\left(\frac{\log(T)}{\delta}\right)$
$\leq 2^{-k+1}C+2|X||A|\log\left(\frac{\log(T)}{\delta}\right)=\theta_{k},$
with probability at least $1-\delta$, so $\mathbb{P}\left(\bigcup_{k\in[M]}\{\sum_{t\in[T]}C_{t}d_{t,k}h_{t,k}>\theta_{k}\}\right)\leq\delta$. Moreover, notice that with probability at least $1-p-2\delta$ thanks to the definition of corruption robust and Azuma-Hoeffding inequality, it holds simultaneously for all $k$:
$\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})d_{t,k}$
$=\frac{1}{2^{-k-1}}\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})2^{-k-1}d_{t,k}$
$=\frac{1}{2^{-k-1}}\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})d_{t,k}\mathbb{E}[h_{t,k}\mid d_{t,k}]$
$=\frac{1}{2^{-k-1}}\left(\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})d_{t,k}\left(\mathbb{E}[h_{t,k}\mid d_{t,k}]-h_{t,k}\right)+\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})d_{t,k}h_{t,k}\right)$
$\leq\frac{1}{2^{-k-1}}\left(L\sqrt{2\ln\left(\frac{\log(T)}{\delta}\right)\sum_{t\in[T]}d_{t,k}}+\sqrt{\beta_{1}\sum_{t\in[T]}d_{t,k}}+(\beta_{2}+\beta_{3}\theta_{k})\max_{t\in[T]}d_{t,k}\right)$
$\leq\mathcal{O}\left(\frac{1}{2^{-k-1}}\left(\left(\sqrt{\beta_{1}}+L\sqrt{\log\left(\frac{T}{\delta}\right)}\right)\sqrt{T}\max_{t\in[T]}d_{t,k}+(\beta_{2}+\beta_{3}\theta)\max_{t\in[T]}d_{t,k}\right)\right),$
noticing that $\mathbb{E}\left[d_{t,k}\left(\mathbb{E}\left[h_{t,k}|d_{t,k}\right]-h_{t,k}\right)\right]=\mathbb{E}\left[h_{t,k}|d_{t,k}\right]-\mathbb{E}[h_{t,k}]d_{t,k}=\mathbb{E}\left[h_{t,k}|d_{t,k}\right]-\mathbb{E}\left[h_{t,k}|d_{t,k}\right]=0$, since the expectation is taken w.r.t. the randomization of Algorithm 4 and the distribution generated given the external probability of receiving feedback $w_{t}$.
To conclude with probability at least $1-p-2\delta$:
$\sum_{t\in[T]}$ $\overline{r}^{\top}(q^{*}-q_{t})\mathbb{I}\left(w_{t}\geq\frac{1}{T}\right)$
$\leq\sum_{k\in[M]}\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t})d_{t,k}$
$\leq\mathcal{O}\left(\sqrt{\beta_{1}T}\max_{t\in[T]}\frac{1}{w_{t}}+(\beta_{2}+\beta_{3}|X||A|\log(\nicefrac{{\log(T)}}{{\delta}}))\max_{t\in[T]}\frac{1}{w_{t}}+\beta_{3}\log(T)C\right)$
$\leq\mathcal{O}\left(\left(\sqrt{\beta_{1}^{\prime}T}+\beta_{2}^{\prime}\right)\nu_{T}+\beta_{3}^{\prime}C\right),$
with $\sqrt{\beta_{1}}\geq\mathcal{O}(L\sqrt{\log(\nicefrac{{T}}{{\delta}})})$. Notice that the analogous reasoning can be applied to the positive cumulative constraints violation with parameters $\beta_{4},\beta_{5},\beta_{6}$. ∎
Algorithm 4 Adapted STABILIZE Jin et al. (2024)
1:$C$, $\delta\in(0,1)$
2:Initialize $M=\log(T)$ instance of Algorithm 2, each instance $k\in[M]$ initialized with corruption parameter:
$\theta_{k}:=2^{-k+1}C+2|X||A|\log\left(\frac{\log(T)}{\delta}\right)$
3:for $t\in[T]$ do
4:     Observe $w_{t}$, probability of receiving feedback.
5:     if $w_{t}>\frac{1}{T}$ then
6:         Let $k_{t}$ be such that $w_{t}\in(2^{-k_{t}-1},2^{-k_{t}}]$
7:         Choose $\pi_{t}$ as policy proposed by instance $k_{t}$
8:         If the algorithm receives feedback send it to instance $k_{t}$ with probability $\frac{2^{-k_{t}-1}}{w_{t}}$
9:     if $w_{t}\leq\frac{1}{T}$ then
10:         Propose random policy $\pi_{t}$
Corollary 6.
Being $j^{*}$ such that $C\in(2^{j^{*}-1},2^{j^{*}}]$ then with probability at least $1-11\delta$ it holds:
$\max_{i\in[m]}\sum_{t\in[T]}\left[\mathbb{E}[g_{t,i}]^{\top}q_{t}^{j^{*}}-\alpha_{i}\right]^{+}\leq\sqrt{\beta_{4}T}\nu_{T,j^{*}}+\beta_{5}\nu_{T,j^{*}}+2\beta_{6}C,$
with $\sqrt{\beta_{4}}=\mathcal{O}\left(L|X|\sqrt{|A|\ln(\nicefrac{{mT|X||A|}}{{\delta}})}\right)$, $\beta_{5}=\mathcal{O}\left(|X|^{2}|A|^{2}\log(T)\log\left(\nicefrac{{\log(T)}}{{\delta}}\right)\right)$and $\beta_{6}=\mathcal{O}\left(\ln(T)^{2}|X||A|\right)$
Corollary 7.
Being $j^{*}$ such that $C\in(2^{j^{*}-1},2^{j^{*}}]$ then with probability at least $1-11\delta$ it holds:
$\sum_{t\in[T]}\overline{r}^{\top}(q^{*}-q_{t}^{j^{*}})\leq\sqrt{\beta_{1}T}\nu_{T,j^{*}}+\beta_{2}\nu_{T,j^{*}}+2\beta_{3}C,$
where $\sqrt{\beta_{1}}=\mathcal{O}\left(L|X|\sqrt{|A|\ln(\nicefrac{{T|X||A|}}{{\delta}})}\right)$, $\beta_{2}=\mathcal{O}\left(|X|^{2}|A|^{2}\log(T)\log\left(\nicefrac{{\log(T)}}{{\delta}}\right)\right)$ and $\beta_{3}=\mathcal{O}\left(\ln(T)^{2}|X||A|\right)$